OpenStudy (anonymous):

can somebody please explain this unit: Kg.sec^-1

OpenStudy (anonymous):

Kilogram per second

OpenStudy (anonymous):

ok, why is the second to the ^-1? also if I was to convert this to N per second what would i do?

OpenStudy (anonymous):

multiply by some N/kilogram

OpenStudy (anonymous):

so multiply by 9.81/1000

OpenStudy (anonymous):

do u mean N to be Newton

OpenStudy (anonymous):

yes Newtons per second

OpenStudy (anonymous):

then just multiply by 9.81 (the value of acceleration due to gravity take care of its unit )

OpenStudy (anonymous):

Oh!, no M...

OpenStudy (anonymous):

Flow rate

OpenStudy (anonymous):

so i would have to: say i had 1000 kg.sec^-1 then i would: 1000*9.81/1000 = 9.81 N/s is that right?

OpenStudy (anonymous):

first check is it kg-wt mentioned.. if then no need to divide by 1000

OpenStudy (anonymous):

see 1kgwt =1*9.81 N = 9.81N

OpenStudy (anonymous):

what i have is: J = 1000 Kg.sec^-1 need to present it in Newtons per sec to match the rest of my question, the rest is either in newtons or needs to be converted to newtons.

OpenStudy (anonymous):

ok but unless mentioned kg-wt u cannot mulltiply with g

OpenStudy (anonymous):

is it related to work-energy problem

OpenStudy (anonymous):

mechanical its for dampners

OpenStudy (anonymous):

damping coefficient

OpenStudy (anonymous):

then u need to check the co-efficient an further if u have any quantity whose unit is Newton /kg u may multiply with that...

OpenStudy (anonymous):

ok thanks you have helped greatly, very much apprciated.

OpenStudy (anonymous):

welcome but i am not sure how far it helped u

OpenStudy (anonymous):

also it says, J can equal: 1000 N.sec.m^1 what is that?